# What Is That Talk About? A Video-to-Text Summarization Dataset for Scientific Presentations

### v3

@@ Abstract
Transforming recorded videos into concise and accurate textual summaries is a growing challenge in multimodal learning. This paper introduces VISTA, a dataset specifically designed for video-to-text summarization in scientific domains. VISTA contains 18,599 recorded AI conference presentations paired with their corresponding paper abstracts. We benchmark the performance of state-of-the-art large models and apply a plan-based framework to better capture the structured nature of abstracts. Both human and automated evaluations confirm that explicit planning enhances summary quality and factual consistency. However, a considerable gap remains between models and human performance, highlighting the challenges of scientific video summarization. [^1] Code and dataset are available here.
@@ 1 Introduction
Large multimodal models (LMMs), which integrate components from different modalities through cross-modal alignment training Koh et al. (2023); Cheng et al. (2023); Li et al. (2024a); Ahn et al. (2024); Fu et al. (2025); Wu et al. (2025), have achieved considerable progress in video-to-text summarization tasks for general-purpose content such as YouTube, movies, and news videos Li et al. (2020); Lin et al. (2023); Krubiński and Pecina (2023); Hua et al. (2024); Chen et al. (2024a); Zhang et al. (2024a); Qiu et al. (2024); Patil et al. (2024); Mahon and Lapata (2024a, b). However, many recent studies have highlighted that these LMMs exhibit reduced performance in scientific contexts, particularly when processing technical terminology and scientific visual elements like figures and tables Li et al. (2024c); Lu et al. (2024); Yue et al. (2024); Hu et al. (2024a); Bai et al. (2024); Liang et al. (2024); Patil et al. (2024); Huang et al. (2024). This performance gap might be largely attributed to the absence of specialized training datasets for multimodal scientific content Chen et al. (2024c); Hu et al. (2024b); Pramanick et al. (2024); Zhang et al. (2024b).
@@
Figure 1: An example from VISTA: a video paired with its abstract. The paper Mallen et al. (2023) was presented at ACL 2023 and received the Best Video Recordings award.
@@
Thus, we introduce VISTA (Video to Scientific Abstract), an English dataset for video-to-text summarization in scientific domains. VISTA consists of 18,599 aligned pairs of conference presentation recordings and their corresponding paper abstracts, collected from leading conferences in computational linguistics (ACL Anthology including ACL, EMNLP, NAACL, EACL, Findings of *ACL) and machine learning (ICML and NeurIPS). Figure 1 illustrates an example selected from VISTA: a conference presentation video (top) paired with the abstract of the corresponding paper (bottom).
@@
We benchmark VISTA using several state-of-the-art (SOTA) large models, including closed-source LMMs (Claude 3.5 Sonnet, Gemini 2.0, GPT-o1), as well as video-specific open-source LMMs (Video-LLaMA, Video-ChatGPT, mPLUG-Owl3, etc.; Zhang et al., 2023; Maaz et al., 2024; Lin et al., 2024a; Ye et al., 2024; Li et al., 2024b, 2025). For comparison, we also include strong baselines: the text-to-text LLaMA-3.1 Touvron et al. (2023) and the audio-to-text Qwen2-Audio Chu et al. (2024). Experiments across zero-shot, QLoRA, and full fine-tuning settings reveal that in-domain fine-tuning improves summarization performance across different large models, and video-based models generally outperform text- and audio-based models on our dataset. However, simpler end-to-end approaches may often struggle to capture the underlying structure of scientific abstracts.
@@
To address this, we explore a plan-based approach, which has been shown to improve coherence and factual grounding through a predefined planning component Liu and Chen (2021); Narayan et al. (2021, 2023). Unlike direct end-to-end generation, plan-based method could leverage the fact that scientific abstracts often follow a well-defined format Takeshita et al. (2024). By explicitly modeling the latent structure of the abstract through a sequence of intermediate plan questions, the summary generation process is better guided. Empirical results confirm that the plan-based method outperforms existing SOTA models in terms of summary quality and factual accuracy. Nevertheless, despite these improvements, all candidate models still struggle with hallucinations and factual errors.
@@
  • We present VISTA, a novel large-scale multimodal dataset with 18,599 video-abstract pairs, tailored for summarizing scientific presentations from video recordings.
@@
  • We apply a plan-based framework that improves upon SOTA video LMMs on summary quality and factual accuracy.
@@ 2 Related Work
generates coherent summaries by integrating multimodal information Hua et al. (2024), supported by datasets like MSS Li et al. (2017), VideoXum Lin et al. (2024b), MMSum Qiu et al. (2024), Hierarchical3D Papalampidi and Lapata (2023), and LfVS-T Argaw et al. (2024), spanning tasks from instructional videos to general web content Li et al. (2017); Zhou et al. (2018); Li et al. (2019, 2020); Liu and Wan (2021); Fu et al. (2021); Krubiński and Pecina (2023); Han et al. (2023); He et al. (2023); Hua et al. (2024); Islam et al. (2024); Qiu et al. (2024). Technical advancements include hierarchical attention models Sanabria et al. (2018), extractive methods using multimodal features Cho et al. (2021); Krubiński and Pecina (2023), and hybrid extractive-abstractive frameworks Ramakrishnan and Ngan (2022); Papalampidi and Lapata (2023). Transformer-based systems have further improved performance Krubiński and Pecina (2023); Li et al. (2020); Shang et al. (2021); Mahon and Lapata (2024a). However, challenges in summarizing academic videos remain under-explored.
@@
condenses complex scholarly content into concise formats Cachola et al. (2020); Ju et al. (2021); Sotudeh and Goharian (2022); Liu and Demberg (2023), supported by datasets like TalkSumm Lev et al. (2019) for academic video transcripts, SumSurvey Liu et al. (2024b) for survey papers, ACLSum Takeshita et al. (2024) for ACL discourse, and SciNews Liu et al. (2024a) for simplifying research for broader audiences. M3AV Chen et al. (2024c) supports tasks like ASR, TTS, and slide-script generation. Methods like HAESum Zhao et al. (2024) and SAPGraph Qi et al. (2022) improve discourse and structural summarization, while CiteSum Mao et al. (2022) and SSR Fatima and Strube (2023) focus on scalability and audience-specific customization. Despite these efforts, scientific summarization remains a challenging domain due to the inherent complexity and diversity of scholarly texts.
@@
employs structured representations to improve summary quality and reduce hallucinations Narayan et al. (2021); Amplayo et al. (2021); Wang et al. (2022); Narayan et al. (2023). Research focuses on text-only planning with elements like entities Narayan et al. (2021); Liu and Chen (2021); Huot et al. (2024), keyword prompts Creo et al. (2023), and question-answer pairs Narayan et al. (2023). Examples include PlanVerb Canal et al. (2022), which converts task plans into natural language via semantic tagging, and domain-specific approaches in dialogue summarization that align with knowledge structures for improved quality Srivastava et al. (2024). Blueprint-based frameworks utilize intermediate plans such as question-answer pairs to create coherent narratives for visual storytelling Liu et al. (2023). However, plan-based strategies for multimodal tasks, particularly video-to-text summarization, have received limited attention.
3 VISTA Dataset
@@
VISTA is derived from computational linguistics and machine learning conferences, including ACL Anthology (ACL, EMNLP, NAACL, EACL, Findings of *ACL), ICML, and NeurIPS, covering content from 2020 to 2024. All materials (paper abstracts and video recordings) are contributed by the respective paper authors, ensuring narrative consistency. Since these metadata are stored in XML/JSON files on their respective websites, no further preprocessing (e.g. extracting abstracts from PDFs) is required. We collect paper titles, author lists, paper abstracts, links to papers, and presentation videos, in accordance with platform terms for academic research purposes (or obtain written confirmation). [^2] We discuss copyright in Appendix A. To maintain one-to-one video-to-text alignments, we exclude samples that may cover multiple papers (e.g., tutorials, invited talks) and videos shorter than one minute or longer than 30 minutes.
@@
The data are sourced directly from official proceedings websites, including textual summaries and presentation videos authored/recorded by corresponding researchers, eliminating the need for additional annotations. We verify the data quality through both human and automated checks. We discuss quality control guidelines and results in Appendix Figure 10 and Appendix B, respectively.
@@
Dataset Language Domain #Videos VideoLen SummaryLen
@@
Shot2Story20K (Han et al., 2023) English Open 20.0K 0.3 201.8
@@
Table 1: Comprison of video-to-text summarization datasets. #Videos $=$ the number of videos, whereas VideoLen and SummaryLen refer to the mean video duration (in minutes) and the average number of summary tokens.
@@
Dataset Comparison and Statistics
Table 1 compares VISTA with several existing video-to-text summarization datasets. While many focus on open-domain (e.g., MMSum, Instruct-V2Xum) or focus on specific areas like news (MLASK, MM-AVS) and activities (VideoXum), VISTA is tailored for summarizing scientific presentations, addressing a distinct niche in video-to-text summarization. On average, it features longer inputs (6.8 minutes) than VideoXum (2.1 minutes) and MSS (3.4 minutes), as well as longer summaries (192.6 tokens), compared to YouCook2 (67.8 tokens) and VideoXum (49.9 tokens).
Table 2 summarizes the dataset statistics: videos average 6.76 minutes and 16.36 shots (we use PySceneDetect with ContentDetector to calculate video shots), while summaries contain 192.62 tokens on average across 7.19 sentences. The average dependency tree depth (Avg. Depth of Dep Tree) is 6.02, indicating the syntactic complexity of the summaries. Meanwhile, the Type-Token Ratio (TTR) is 0.62, reflecting lexical diversity. Both metrics are calculated using spaCy. Diversity metrics Li et al. (2016), which measure the variety of unique n-grams, yield Distinct-1, Distinct-2, and Distinct-3 scores of 0.62, 0.93, and 0.97, respectively. Figure 3 visualizes key attributes: most summaries remain under 250 tokens and 10 sentences, and most videos last fewer than 10 minutes with under 30 shots. In Appendix C, we present a random sample from the VISTA dataset.
Figure 4: GPT-o1 generates plans based on reference summaries. Each question $q_{i}$ corresponds to summary sentence $t_{i}$ which we assume constitutes its answer. Index $i$ ranges from $1$ to the number of summary sentences.
@@ 4 Benchmarking VISTA
We formalize the task of summarizing recorded scientific videos as follows: Let $v$ and $s$ denote a video (or its transcript/audio) and its paired summary from dataset $D=\{(v_{1},s_{1}),(v_{2},s_{2}),\ldots,(v_{n},s_{n})\}$, where $n$ signifies the number of video-abstract pairs. The objective is to train a (multimodal) model $\mathcal{M}$ to learn the conditional probability distribution $P(s\leavevmode\nobreak\ |\leavevmode\nobreak\ v)$. Given a new video, the trained model $\mathcal{M}$ is expected to generate an appropriate summary.
A challenge in video-to-text summarization is structuring the generated summaries in a coherent and faithful manner. Directly learning the mapping from $v$ to $s$ could lead to inadequate outputs, as the model lacks explicit guidance on how to organize and present the extracted information. Scientific abstracts often follow a relatively well-defined structure, making them suitable for a more structured generation approach Takeshita et al. (2024). We follow previous work Narayan et al. (2021); Liu et al. (2023); Narayan et al. (2023) in adopting a plan-based framework that introduces an intermediate representation to capture latent structure more effectively than simpler end-to-end approaches. Specifically, given input video $v$, we first generate plan $p$, which consists of a sequence of automatically generated questions $\{q_{1},q_{2},\ldots,q_{m}\}$, each corresponding to a sentence to be verbalized in the summary. The plan explicitly controls the structure of the summary as a whole and the content of each of its sentences (which are meant to answer the questions in the plan). The model is then trained to learn the extended conditional probability distribution $P(s\leavevmode\nobreak\ |\leavevmode\nobreak\ v,p)$, ensuring that the generated summaries follow the structure and flow of plan $p$.
@@
We hypothesize that summary sentences can be viewed as responses to plan questions directly associated with them. This idea is inspired by the theory of Questions Under Discussion (QUD) Roberts (2012); Wu et al. (2023b); Suvarna et al. (2024), which posits that discourse often revolves around a set of questions that guide the structure and interpretation of the conversation.
We leverage GPT-o1 Achiam et al. (2023) to generate silver-standard plans based on reference summary sentences and their preceding context. As shown in Figure 4, for example, question $q_{3}$ is generated based on target sentence $t_{3}$ and the summary sentences preceding it (i.e., $t_{1}$ and $t_{2}$), and so on. As a result, the question sequence preserves the order of sentences in the reference summaries, ensuring that the plan maintains a natural and coherent flow consistent with the structure of reference summaries. The prompt used to generate plan questions is provided in Appendix Figure 12.
@@
We train two independent models corresponding to Plan Generation (PG) and Summary Generation (SG). The PG module is trained on pairs of $(v,p)$ samples, where $v$ represents the input and $p$ is the silver-standard plan. The SG module is trained on tuples $([v;p],s)$, where $[v;p]$ is the concatenation of the input $v$ and its plan $p$. During inference, the trained PG module predicts plan $\hat{p}$ for input $v$, and the tuple $[v;\hat{p}]$ is fed into the SG module to generate the final summary. Both modules have the same backbone but are trained independently.
Method Model R1 R2 RLsum SacreBLEU Meteor BERTscore CIDEr-D VideoScore FactVC
Zero-shot Learning Claude 3.5 Sonnet 27.71 5.59 24.14 3.14 17.53 82.57 1.32 1.91 50.11
Gemini 2.0 27.82 5.66 24.29 4.22 17.83 82.64 1.47 2.02 52.02
GPT-o1 27.90 5.69 24.37 4.38 17.90 82.63 1.61 2.17 51.36
[…]
@@
Table 3: Model performance on VISTA dataset. In Plan-mPlug-Owl3 , only the Plan Generation (PG) module is trained. Plans generated by the PG module on the test set serve as input to the Summary Generation (SG) module for zero-shot inference (no training is applied to the SG module). Symbols  and indicate that the performance of Plan-mPlug-Owl3 is significantly ($p<0.05$) different from LLaVA-NeXT-Interleave (third best) and mPLUG-Owl3 (second best), when using a paired t-test.

### v4

@@ Abstract
Transforming recorded videos into concise and accurate textual summaries is a growing challenge in multimodal learning. This paper introduces VISTA, a dataset specifically designed for video-to-text summarization in scientific domains. VISTA contains 18,599 recorded AI conference presentations paired with their corresponding paper abstracts. We benchmark the performance of state-of-the-art large models and apply a plan-based framework to better capture the structured nature of abstracts. Both human and automated evaluations confirm that explicit planning enhances summary quality and factual consistency. However, a considerable gap remains between models and human performance, highlighting the challenges of our dataset. This study aims to pave the way for future research on scientific video-to-text summarization. The project information is available at https://dongqi.me/projects/VISTA.
@@ 1 Introduction
Large multimodal models (LMMs), which integrate components from different modalities through cross-modal alignment training Koh et al. (2023); Cheng et al. (2023); Li et al. (2024a); Ahn et al. (2024); Fu et al. (2025); Wu et al. (2025), have achieved considerable progress in video-to-text summarization tasks for general-purpose content such as YouTube, movies, and news videos Li et al. (2020); Lin et al. (2023); Krubiński and Pecina (2023); Hua et al. (2024); Chen et al. (2024a); Zhang et al. (2024a); Qiu et al. (2024); Patil et al. (2024); Mahon and Lapata (2024a, b). However, many recent studies have highlighted that these LMMs exhibit reduced performance in scientific contexts, particularly when processing technical terminology and scientific visual elements like figures and tables Li et al. (2024b); Lu et al. (2024); Yue et al. (2024); Bai et al. (2024); Liang et al. (2024); Patil et al. (2024); Huang et al. (2024). This performance gap might be largely attributed to the absence of specialized datasets for multimodal scientific content Chen et al. (2024c); Hu et al. (2024); Pramanick et al. (2024); Zhang et al. (2024b).
@@
Figure 1: An example from VISTA: a conference presentation video (top) paired with the abstract of the corresponding paper (bottom). This data sample Mallen et al. (2023) was presented at ACL 2023 and received the Best Video Recordings award.
@@
Thus, we introduce VISTA (Video to Scientific Abstract), an English dataset for video-to-text summarization in scientific domains. VISTA consists of 18,599 aligned pairs of conference presentation recordings and their corresponding paper abstracts, collected from leading conferences in computational linguistics (ACL Anthology including ACL, EMNLP, NAACL, EACL, Findings of *ACL) and machine learning (ICML and NeurIPS). Figure 1 illustrates an example selected from VISTA.
@@
We use the abstract of the paper as a proxy for the summary of the video and benchmark VISTA using several state-of-the-art (SOTA) large models, including closed-source LMMs (Claude 3.5 Sonnet, Gemini 2.0, GPT-o1), as well as video-specific open-source LMMs (Video-LLaMA, Video-ChatGPT, mPLUG-Owl3, etc.; Zhang et al., 2023; Maaz et al., 2024; Lin et al., 2024a; Ye et al., 2025; Li et al., 2025, 2024c). For comparison, we also include strong baselines: text-to-text model LLaMA-3.1 Touvron et al. (2023) and audio-to-text model Qwen2-Audio Chu et al. (2024). Experiments across zero-shot, QLoRA, and full fine-tuning settings reveal that in-domain fine-tuning improves summarization performance across different large models, and video-based models generally outperform text- and audio-based models on our dataset. However, end-to-end approaches may often struggle to capture the underlying structure of scientific abstracts Liu et al. (2025).
@@
To address this, we explore a plan-based approach, which has been shown to improve coherence and factual grounding through a predefined planning component Narayan et al. (2021, 2023); Liu et al. (2025). Unlike direct end-to-end generation, plan-based methods can leverage the fact that scientific abstracts often follow a well-defined format Takeshita et al. (2024). By explicitly modeling the latent structure of the summary through a sequence of intermediate plans, the summary generation process can be better guided. Empirical results confirm that the plan-based method outperforms existing SOTA models in terms of summary quality and factual accuracy. This work also lays the groundwork for future investigations into the multimodal summarization of scientific videos.
@@
  • We present VISTA, a novel large-scale multimodal dataset with 18,599 video-summary pairs, tailored for summarizing scientific presentations from video recordings.
@@
  • We leverage a plan-based approach that consistently improves summary quality and factual accuracy over SOTA models.
@@ 2 Related Work
generates coherent summaries by integrating multimodal information Hua et al. (2024), supported by datasets like MSS Li et al. (2017), VideoXum Lin et al. (2024b), MMSum Qiu et al. (2024), Hierarchical3D Papalampidi and Lapata (2023), and LfVS-T Argaw et al. (2024), spanning tasks from instructional videos to general web content Li et al. (2017); Zhou et al. (2018); Li et al. (2019, 2020); Liu and Wan (2021); Fu et al. (2021); Liu et al. (2022); Krubiński and Pecina (2023); Han et al. (2025); He et al. (2023); Hua et al. (2024); Islam et al. (2024); Qiu et al. (2024). Technical advancements include hierarchical attention models Sanabria et al. (2018), extractive methods using multimodal features Cho et al. (2021); Krubiński and Pecina (2023), and hybrid extractive-abstractive frameworks Ramakrishnan and Ngan (2022); Papalampidi and Lapata (2023). Transformer-based systems have further improved performance Krubiński and Pecina (2023); Li et al. (2020); Shang et al. (2021); Mahon and Lapata (2024a). However, challenges in summarizing academic videos remain under-explored.
@@
condenses complex scholarly content into concise formats Cachola et al. (2020); Ju et al. (2021); Liu et al. (2023b); Liu and Demberg (2023), supported by datasets like TalkSumm Lev et al. (2019) for academic video transcripts, SumSurvey Liu et al. (2024b) for survey papers, ACLSum Takeshita et al. (2024) for ACL discourse, and SciNews Liu et al. (2024a) for simplifying research for broader audiences. M3AV Chen et al. (2024c) supports tasks like ASR, TTS, and slide-script generation. Methods like RST-LoRA Liu and Demberg (2024) and RSTformer Liu et al. (2023b) improve discourse and structural summarization, while CiteSum Mao et al. (2022) and SSR Fatima and Strube (2023) focus on scalability and audience-specific customization. Despite these efforts, scientific summarization remains a challenging domain due to the inherent complexity and diversity of scholarly texts.
@@
employs structured representations to improve summary quality and reduce hallucinations Narayan et al. (2021); Amplayo et al. (2021); Wang et al. (2022); Narayan et al. (2023); Liu et al. (2025). Research focuses on text-based planning with elements like entities Narayan et al. (2021); Liu and Chen (2021); Huot et al. (2024), keyword prompts Creo et al. (2023), and question-answer pairs Narayan et al. (2023). Examples include PlanVerb Canal et al. (2022), which converts task plans into natural language via semantic tagging, and domain-specific approaches that align with knowledge structures for improved quality Srivastava et al. (2024). Blueprint-based frameworks utilize intermediate plans to create coherent narratives for visual storytelling Liu et al. (2023a). However, plan-based strategies for multimodal tasks, particularly video-to-text summarization, have received limited attention.
3 The VISTA Dataset
@@
VISTA is derived from computational linguistics and machine learning conferences, including ACL Anthology (ACL, EMNLP, NAACL, EACL, Findings of *ACL), ICML, and NeurIPS, covering content from 2020 to 2024. All materials (paper abstracts and video recordings) are contributed by the respective paper authors, ensuring narrative consistency. Since these metadata are stored in XML/JSON files on their respective websites, no further data preprocessing (e.g., extracting abstracts from PDFs) is required. We collect paper titles, author lists, paper abstracts, links to papers, and presentation videos, in accordance with platform terms for academic research purposes (or obtain written confirmation). [^1] We discuss copyright in Appendix A. To maintain one-to-one video-to-text alignments, we exclude samples that may cover multiple papers (e.g., tutorials, invited talks) and videos shorter than one minute or longer than 30 minutes.
@@
We verify the data quality through both manual and automated checks. We discuss quality control guidelines and the results in Appendix Figure 10 and Appendix B, respectively.
@@
Dataset Language Domain #Videos VideoLen SumLen
@@
Shot2Story20K (Han et al., 2025) English Open 20.0K 0.3 201.8
@@
Table 1: Comprison of video-to-text summarization datasets. #Videos $=$ the number of videos, whereas VideoLen and SumLen refer to the average of video duration (in minutes) and the average number of summary tokens.
@@
Dataset Comparison and Statistics
Table 1 compares VISTA with several existing video-to-text summarization datasets. While many focus on open-domain (e.g., MMSum, Instruct-V2Xum) or areas like news (MLASK, MM-AVS) and activities (VideoXum), VISTA is tailored for summarizing scientific presentations. On average, it features longer inputs (6.8 minutes) than VideoXum (2.1 minutes) and MSS (3.4 minutes), as well as longer summaries (192.6 tokens), compared to YouCook2 (67.8 tokens) and VideoXum (49.9 tokens).
@@
Table 2 summarizes the VISTA dataset statistics: Videos average 6.76 minutes and 16.36 shots (we use PySceneDetect with ContentDetector to calculate video shots), while summaries contain 192.62 tokens on average across 7.19 sentences. The average dependency tree depth (Avg. Depth of Dep Tree) is 6.02, indicating the syntactic complexity of the summaries. Meanwhile, the Type-Token Ratio (TTR) is 0.62, reflecting lexical diversity. Both metrics are calculated using spaCy. Diversity metrics Li et al. (2016), which measure the variety of unique n-grams, yield Distinct-1, Distinct-2, and Distinct-3 scores of 0.62, 0.93, and 0.97, respectively. Figure 3 visualizes key attributes: Most summaries remain under 250 tokens and 10 sentences, and most videos last fewer than 10 minutes with under 30 shots. In Appendix C, we present a random sample from the VISTA dataset.
Figure 4: GPT-o1 generates plans based on reference summaries. Each question $q_{i}$ corresponds to a summary sentence $t_{i}$, which we assume constitutes its answer. Index $i$ ranges from $1$ to the number of summary sentences.
@@ 4 Benchmarking VISTA
We formalize the task of summarizing recorded scientific videos as follows: Let $v$ and $s$ denote a video (or its transcript/audio) and its paired summary from dataset $D=\{(v_{1},s_{1}),(v_{2},s_{2}),\ldots,(v_{n},s_{n})\}$, where $n$ signifies the number of video-summary pairs. The objective is to train a (multimodal) model $\mathcal{M}$ to learn the conditional probability distribution $P(s\leavevmode\nobreak\ |\leavevmode\nobreak\ v)$. Given a new video, the trained model $\mathcal{M}$ is expected to generate an appropriate summary.
A challenge in video-to-text summarization is structuring the generated summaries in a coherent and faithful manner. Directly learning the mapping from $v$ to $s$ could lead to inadequate outputs, as the model lacks explicit guidance on how to organize and present the extracted information Mahon and Lapata (2024a). Scientific abstracts often follow a relatively well-defined structure, making them suitable for a more structured generation approach Takeshita et al. (2024). We follow previous work Narayan et al. (2021, 2023) in adopting a plan-based framework that introduces an intermediate representation to capture latent structure more effectively than simpler end-to-end approaches. Specifically, given input $v$, we first generate a plan $p$, which consists of a sequence of automatically generated questions $\{q_{1},q_{2},\ldots,q_{m}\}$, each corresponding to a sentence to be verbalized in the summary. The plan explicitly controls the structure of the summary as a whole and the content of each of its sentences (which are meant to answer the questions in the plan). The model is then trained to learn the extended conditional probability distribution $P(s\leavevmode\nobreak\ |\leavevmode\nobreak\ v,p)$, ensuring that the generated summaries follow the structure and flow of plan $p$.
@@
We hypothesize that summary sentences can be viewed as responses to plan questions, where the plan consists of an ordered sequence of questions directly associated with the target content. This idea is inspired by the theory of Question Under Discussion (QUD; Roberts (2012); Wu et al. (2023b); Suvarna et al. (2024)), which posits that discourse often revolves around a set of questions that guide the structure and interpretation of the conversation.
We leverage GPT-o1 Achiam et al. (2023) to generate silver-standard plans based on reference summary sentences and their preceding context. As shown in Figure 4, for example, question $q_{3}$ is generated based on target sentence $t_{3}$ and the summary sentences preceding it (i.e., $t_{1}$ and $t_{2}$), and so on. As a result, the question sequence preserves the order of sentences in the reference summaries, ensuring that the plan maintains a natural and coherent flow consistent with the structure of reference summaries. The prompt used to generate plan questions is provided in Appendix Figure 12. We discuss the quality of the silver-standard plans through manual investigation in Appendix G.
@@
We train two independent modules corresponding to Plan Generation (PG) and Summary Generation (SG). The PG module is trained on pairs of $(v,p)$ samples. The SG module is trained on tuples $([v;p],s)$, where $[v;p]$ is the concatenation of the input $v$ and its plan $p$. During inference, the trained PG module predicts plan $\hat{p}$ for input $v$, and the tuple $[v;\hat{p}]$ is fed into the SG module to generate the final summary. Both modules have the same backbone but are trained independently.
Method Model Open-source R1 R2 RLsum SacreBLEU Meteor BERTscore CIDEr-D VideoScore FactVC
Zero-shot Learning LLaMA-3.1transcript 23.68 4.22 21.39 2.70 14.62 80.93 1.17 1.53 34.32
@@
Table 3: Model performance on VISTA dataset. In Plan-mPlug-Owl3, only the PG module is trained. Plans generated by the PG on the test set serve as input to the SG module for zero-shot inference (no training is applied to the SG module). Symbols  and indicate that the performance of Plan-mPlug-Owl3 is significantly ($p<0.05$) different from LLaVA-NeXT-Interleave (third best) and mPLUG-Owl3 (second best), when using the paired t-test.